RC meeting and Run Plan, Tuesday November 22nd

From SBS wiki
Jump to: navigation, search

Return to the RC Meeting Calendar

RC Meeting Information

Time:at 3:30pm today

  • Location: Counting house conference room

RC Updates

Summary of swing 11/21 - day 11/22

  • Maintenace in South linac scroll/turbo pumps and then 2L17 HV issues [opportunistic EPR calibration completed around 5 pm] yesterday
  • Beam checkout completed around 10 pm yesterday, and then continued production run on polarized He3 target @45uA beam current with prescale: ps3=0, ps5=3 and all other -1
  • TO reported the issue of fiber temp above alarm limits around 6:30 am today
  • Arun and Bill tried to fix target issue remotely
  • TO did NMR while Bill was working on oven temperature with his approval
  • Target Polarization was only 1.6% after NMR, so contacted Bill again
  • Bill and Arun believed that there is some issue with target, so decided to go to controlled access (around 10 am)
  • RadCon assisted to have preliminary assessment https://logbooks.jlab.org/entry/4089408
  • We are still controlled access and shift crews (day and swing swift) are standby

Short term run plan

  • Access the conditions inside the target enclosure, and install another cell
  • Integration of SBS GEMS to Main DAQ
  • GEMs luminosity scan
  • Change in coincidence window size
  • Take run with lower BBCAL threshold (~1 GeV) to accept pions

Issues

  • Shift schedule unfilled for next week
  • Dump right Ion chamber cabling needs to be changed (Jerry Kowal is aware of it)
  • Two channels of hodoscope have issue https://logbooks.jlab.org/entry/4087074

Hall Access Request

  • William Henry (target)
  • Sean Jeffas (GEMs)
  • Jiwan Poudel (trigger)
  • Anuruddha Rathnayake (GEMs)
  • Carlos Ayerbe (GRICH)

Summary of GEn data taking

  • 58 out of 141 days (41.13%)
  • Uncorrected accumulated charge on He3: 34.75C out of 232C (14.98%)

Figure: Charge accumulated over time with He3 target